0.000675 M calcium hydroxide


1
Expert's answer
2022-06-09T21:48:03-0400

CM [Ca(OH)2] = 0,000675 M = 6,75 * 10-4

pH = ?


1.

We find the OH– ion concentration

Ca(OH)2 => Ca2+ + 2OH–

1mol Ca(OH)2 => 2mol OH–

6,75 * 10-4 M => X

X = 6,75 * 10-4 * 2 / 1 = 1,35 * 10-3 M


2.

We find the pOH:

pOH = -log [OH–]

pOH = -log (1,35 * 10-3) = 3- log (1,35) = 2,87


3.

Using the sum of pH and pOH equal 14 in any solution, we find the pH:

pH + pOH = 14

pH = 14 - pOH = 14 - 2,87 = 11,13


Answer: pH = 11,13